# Bott-integrability of overtwisted contact structures

### v1

@@ 1. Introduction
Recall that a graph manifold is a $3$-manifold that can be cut along tori into Seifert fibred pieces. In [4, Theorem 1.3] we showed that a closed $3$-manifold admits a Bott-integrable contact structure if and only if it is a graph manifold. We also gave examples of non-integrable contact structures on graph manifolds, and examples of $3$-manifolds (the $3$-sphere $S^{3}$, the $3$-torus $T^{3}$, and $S^{1}\times S^{2}$) where every contact structure is Bott integrable.
@@ 2. The critical set
In [4, Theorem 2.2] we proved the following Liouville type theorem: If $f$ is an integral for the Reeb flow, that is, $\mathrm{d}f(R)=0$, then any regular level set of $f$ is a $2$-torus $T^{2}$, and in terms of circle coordinates $x_{1},x_{2}$ on $T^{2}$ and a suitable transverse coordinate $r$, the contact form $\alpha$ is of Lutz type $\alpha=h_{1}(r)\,\mathrm{d}x_{1}+h_{2}(r)\,\mathrm{d}x_{2}$, and the integral $f$ a function of $r$.
For the case when the integral $f$ is a Morse–Bott function, we also established a neighbourhood theorem for isolated elliptic Reeb orbits in the critical set $\operatorname{Crit}(f)$, that is, an orbit where the Bott integral has a minimum or a maximum; see [4, Theorem 2.4]: In terms of suitable coordinates $(\theta;r,\varphi)$ on a neighbourhood $S^{1}\times D^{2}_{\delta}$ of the critical orbit $S^{1}\times\{0\}$, the Bott integral $f$ equals $f(r)=c\pm r^{2}$, and up to an isotopic deformation of the contact form in a neighbourhood of the critical orbit—changing the Reeb flow but not the Bott integral—the contact form is given by $\mathrm{d}\theta+r^{2}\,\mathrm{d}\varphi$. Also, by Gray stability, the contact structure $\xi=\ker\alpha$ remains unchanged, up to a diffeotopy of $M$, under this deformation.
In the present section we extend the arguments for achieving those results in order to prove neighbourhood theorems for tori and Klein bottles contained in $\operatorname{Crit}(f)$. We then reason as in [4, Section 10.2] to remove those surfaces from the critical set by a $C^{\infty}$-small perturbation of the Bott integral $f$, at the cost of introducing an elliptic and a parallel hyperbolic Reeb orbit into the critical set.
@@ 2.1. Neighbourhood theorems for critical surfaces
Statement (i) follows from the generalised Morse–Bott lemma [1, Lemma 1.7]. In order to prove (ii), we slightly modify the argument for proving [4, Theorem 2.2]. There we constructed covering maps $\mathbb{R}^{2}\rightarrow\{r\}\times T^{2}$ by using the flow of $R$ and $Y$. In the present situation, $Y$ vanishes along the critical set $\{r=0\}$. We circumvent this problem by working with the vector field $Y^{\prime}$ defined by
@@
in other words, we replace $\mathrm{d}f$ in the defining equations (1) for $Y$ by $\mathrm{d}f/r=\pm 2\,\mathrm{d}r$ for $r\neq 0$. This is well defined also in $r=0$, since $r$ is a smooth coordinate on $(-\delta,\delta)$. So $Y^{\prime}$ coincides with $Y/r$ for $r\neq 0$ and extends smoothly as a non-vanishing vector field over $r=0$. The remainder of the argument is now as in the proof of [4, Theorem 2.2]. ∎
Next we describe the neighbourhood of a critical Klein bottle $K$. Recall from [5] or [4, Section 10.1] that a closed tubular neighbourhood $\nu K$ of a Klein bottle $K$ embedded in an orientable $3$-manifold, which is the unique non-trivial interval bundle over $K$, can be described by
@@
The contact form $\alpha$ and the Bott integral $f$ lift to this double cover. In this lift, $\{0\}\times T^{2}=\{r=0\}$ is a component of the critical set. Now, we cannot directly quote the previous result, since we need to have a $\mathbb{Z}_{2}$-equivariant version of that neighbourhood theorem, and we want to achieve a more specific Lutz form. So we need to dig a little deeper into the proof of the neighbourhood theorem for elliptic critical orbits [4, Theorem 2.4], which contains the relevant ideas. At each step in the following argument, we deal with a $\mathbb{Z}_{2}$-equivariant isotopy or a $\mathbb{Z}_{2}$-invariant rescaling of the contact form, all of which are supported in a neighbourhood of $\{r=0\}$. So the actual normal form we finally arrive at is defined on a neighbourhood of $K$ defined by a smaller $\delta$ than the one we started with.
@@
In particular, we have $h_{0}(0)=0$. As in Step 1 of the proof of [4, Theorem 2.4], we first deform $h_{0}$, relative to a neighbourhood of {±$\{\pm\delta\}\times T^{2}$, so that it becomes identically zero near $r=0$; this can be done within the class of odd functions. The convex linear interpolation between the old and the deformed $h_{0}$ gives rise to a strict contact isotopy, defined by a flow along the level surfaces $\{r=\text{const.}\}$. }, and stationary on $\{r=0\}$, just like in the proof of [4, Theorem 2.2]; see equation (3) in that proof and the argument following it.
The $\mathbb{Z}_{2}$-invariance of $\widetilde{R}$, which depends on $r$ only, forces the Reeb vector field to be a multiple of $\partial_{t}$ along $\{r=0\}$. This implies that $h_{1}(0)\neq 0$, and by choosing the signs of the coordinates $r,s$ appropriately, we may assume that $h_{1}(0)>0$ and that the ambient orientation is given by $\mathrm{d}r\wedge\mathrm{d}s\wedge\mathrm{d}\theta$. Then as in Step 2 of the proof of [4, Theorem 2.4], we can make a further deformation such that the contact form looks like $\mathrm{d}s+h_{2}(r)\,\mathrm{d}\theta$ (with a modified, but still odd $h_{2}$) near $\{r=0\}$: for this we need to divide the contact form by an even function $\chi(r)$ equal to $h_{1}(r)$ near $r=0$, and identically equal to $1$ outside a small neighbourhood of $r=0$; so the contact structure remains unchanged. This deformation changes the Reeb vector field, keeping it tangent to the tori $\{r=\text{const.}\}$. }. Thanks to the contact condition, $h_{2}$ satisfies $h_{2}^{\prime}<0$.
@@ 2.2. Removing critical surfaces
For the Klein bottle, we explained in [4, Section 10.2] how to perturb the Bott integral $f$ into one that has two isolated critical Reeb orbits, one elliptic and one hyperbolic, in place of the critical Klein bottle.
@@
In the torus case, we modify the contact form $\alpha=h_{1}(r)\,\mathrm{d}x_{1}+h_{2}(r)\,\mathrm{d}x_{2}$ near $\{r=0\}$ in a way described in the proof of [4, Proposition 2.6]. By a small reparametrisation of the curve $r\mapsto(h_{1}(r),h_{2}(r))$ we may assume that $h_{1}(0)$ and $h_{2}(0)$ are rationally dependent. Then, up to an $\mathrm{SL}(2,\mathbb{Z})$-transformation we may assume that $h_{1}(0)>0$ and $h_{2}(0)=0$, which implies $h_{2}^{\prime}(0)<0$. After a small deformation and a further reparametrisation of the curve $r\mapsto(h_{1}(r),h_{2}(r)$, up to a global constant we may finally assume that $h_{1}$ is identically equal to $1$ near $r=0$, and $h_{2}(r)=-r$ near $r=0$. These modifications change the Reeb vector field, keeping it tangent to the tori $\{r=\text{const.}\}$. }, so $f$ is still a Bott integral. Moreover, near $\{r=0\}$ the new Reeb vector field equals $\partial_{x_{1}}$, so the new Bott integral
@@ 3.1. Euler class and critical orbits
For the case that $L_{f}=\operatorname{Crit}(f)$, that is, in the absence of critical surfaces, the following result was stated implicitly in the proof of [4, Lemma 7.2]. Here and throughout this paper, (co-)homology groups are understood to be with $\mathbb{Z}$-coefficients.
@@ 3.2. The link of critical orbits
In [4, Section 3.2] we gave a summary of the results from [1] concerning the (singular) Liouville foliation defined by the level sets of a Bott integral. The key statement is that up to a diffeomorphism preserving the Liouville foliation, any Bott-integrable Reeb flow can be constructed from some simple building blocks, which are products of a surface (with boundary) with $S^{1}$, or Seifert bundles with singular fibres of order $2$, coming from fixed points of an involution on the surface. The Liouville foliation on these building blocks is defined by the level sets of a Morse–Bott function lifted from a Morse–Bott function on the base, with all fixed points of the involution contained in the critical set of that Morse function.
@@
In [4, Section 3.2.3], we wrote that the Morse–Bott functions on the Seifert fibred building blocks are lifted from Morse functions on the base, when in fact we should have written ”Morse–Bott functions on the base”. The neighbourhood $\nu K$ of a critical Klein bottle is a case in point. As described in detail in [5], $\nu K$ may be thought of as the Seifert bundle over a disc with two singular fibres of order $2$. This corresponds to starting with the annulus $[-\delta,\delta]\times S^{1}$ with Morse–Bott function $f(r,\theta)=r^{2}$ and involution $\tau(r,\theta)=(-r,-\theta)$ having two fixed points. One then obtains $\nu K$ with the described Seifert fibration as the mapping torus of $\tau$.
@@ 4. Seifert manifolds
In [4, Theorem 1.8] we showed that every contact structure on a closed Seifert manifold invariant under the fixed-point free $S^{1}$-action defining the Seifert fibration is Bott integrable. In the present section we deal with overtwisted (but not necessarily $S^{1}$-invariant) contact structures on Seifert manifolds.
Our aim will be to find a Bott-integrable contact structure on any given Seifert manifold that is sufficiently ‘rich’ in the following sense. When we speak of a Bott-integrable contact structure on a manifold with boundary, it is understood that the boundary components are tori, and the contact structure is of Lutz type on collar neighbourhoods of those boundary components. In particular, the boundary is a regular level set of the Bott integral, and the Reeb flow is tangent to the boundary. By the sewing lemma [4, Lemma 3.4], the gluing along such boundaries produces again a Bott-integrable contact structure.
@@
By Eliashberg’s classification [2] of overtwisted contact structures we need only show that every homotopy class of tangent $2$-plane fields on $M$ contains a Bott-integrable overtwisted contact structure. Write $\xi_{0}$ for the given $H_{1}$-complete Bott-integrable contact structure on $M$. We may assume that $\xi_{0}$ is overtwisted; if it is not, we create an elliptic critical orbit (together with a hyperbolic one) as in [4, Section 2.3], and then perform a full Lutz twist, which does not change the homotopy class of $\xi_{0}$ as a tangent $2$-plane field.
Now, given any tangent $2$-plane field $\eta$, there is an obstruction $d^{2}(\xi_{0},\eta)\in H^{2}(M)$ for $\xi_{0}$ to be homotopic to $\eta$ over the $2$-skeleton of $M$. This obstruction is antisymmetric and additive; see [3, Section 4.2]:
@@
see [3, Remark 4.3.4]. Thus, in the presence of $2$-torsion in $H^{2}(M)$ the obstruction class is not detected by the Euler class.
@@
Finally, in order to modify $\xi^{u}$ into a Bott-integrable contact structure homotopic to $\eta$ over all of $M$, it suffices to take a contact connected sum as in [4, Section 6] with a suitable overtwisted contact structure on $S^{3}$, all of which are Bott integrable by [4, Theorem 1.9]. ∎
@@ 4.1.1. Open books
As discussed in [4, Section 7], the manifold $\Sigma\times S^{1}$ may be thought of as an open book with page $\Sigma_{0}$ equal to $\Sigma$ with two open discs removed, and monodromy a couple of Dehn twists (one right-handed, one left-handed) parallel to the two boundary components $\partial\Sigma_{0}\setminus\partial\Sigma$ created by the removal of these discs. For the construction of an $H_{1}$-complete Bott-integrable contact structure on $\Sigma\times S^{1}$, we need the following mild extension of [4, Lemma 3.3]. We formulate it for a single oriented, embedded circle $\gamma\subset\Sigma_{0}$, but it works equally well for a finite collection of pairwise disjoint embedded circles.
@@
Now apply the argument from [4, Lemma 3.3], involving compactly supported cohomology, to the surface obtained by cutting open $\Sigma_{0}$ along $\gamma$ in order to find a primitive $\lambda$ of $\omega$ with the desired properties. ∎
Write $\psi\colon\thinspace\Sigma_{0}\rightarrow\Sigma_{0}$ for the diffeomorphism given by the two boundary-parallel Dehn twists. As shown in [4, Section 7.2], an area form $\omega=\mathrm{d}\lambda$ as just constructed gives rise to a Bott-integrable contact form on $\Sigma\times S^{1}$, where the Bott integral on the mapping torus $\Sigma_{0}(\psi)$ can be taken to be one induced from any Morse function on $\Sigma_{0}$ that only depends on the radial coordinate in the collar neighbourhoods of the boundary (and hence is invariant under $\psi$). On the complement $\Sigma_{0}^{\prime}\subset\Sigma_{0}$ of two collars containing the annuli where the Dehn twists are performed, the mapping torus looks like a product $\Sigma_{0}^{\prime}\times S^{1}$ with contact form $\lambda+C\,\mathrm{d}\theta$, where $C$ is some large positive constant.

### v2

@@ 1. Introduction
Recall that a graph manifold is a $3$-manifold that can be cut along tori into Seifert fibred pieces. In [6, Theorem 1.3] we showed that a closed $3$-manifold admits a Bott-integrable contact structure if and only if it is a graph manifold. We also gave examples of non-integrable contact structures on graph manifolds, and examples of $3$-manifolds (the $3$-sphere $S^{3}$, the $3$-torus $T^{3}$, and $S^{1}\times S^{2}$) where every contact structure is Bott integrable.
@@
Remark 1.3.
Here is a brief comment on the dynamical significance of this theorem. By a result of Paternain [18], a Hamiltonian flow (on a $3$-dimensional energy hypersurface) with an integral whose critical set consists of submanifolds has vanishing topological entropy. In particular, this applies to Bott-integrable Reeb flows. Macarini–Schlenk [16], by contrast, have shown the existence of contact structures that force entropy, in the sense that any associated Reeb flow has positive topological entropy. Other results of this type are due to Alves [1] and Alves–Colin–Honda [2]. In all cases, the relevant property of the contact structure that forces entropy also implies tightness of the contact structure.
In [2, Question 1.5], the authors ask whether there exists an overtwisted contact structure in dimension $3$ that forces entropy. As shown by Yau [24], the contact homology of overtwisted contact structures vanishes. This suggests that the orbit structure of overtwisted Reeb flows may not be rich enough to establish forced entropy by the methods of [1].
In combination with the result of Paternain, our Theorem 1.2 gives a criterion for an overtwisted contact structure not to force entropy. As an example, our results in Section 4 show that for Seifert manifolds the answer to [2, Question 1.5] is negative: no overtwisted contact structure on a Seifert manifold forces entropy.
On the other hand, Theorem 1.2 provides examples of overtwisted contact structures on graph manifolds that force any associated Reeb flow at least to be non-integrable (in the Bott sense); see Section 6. Now, one of two things can happen with these examples: Either (i) entropy is forced, providing a positive answer to [2, Question 1.5], or (ii) there is a Reeb flow of zero topological entropy that cannot be approximated by Bott-integrable Reeb flows. Case (ii) would go some way towards answering a question of Katok [15, Problem 1], who wondered about the existence of such approximations by Liouville-integrable systems.
@@ 2. The critical set
In [6, Theorem 2.2] we proved the following Liouville type theorem: If $f$ is an integral for the Reeb flow, that is, $\mathrm{d}f(R)=0$, then any regular level set of $f$ is a $2$-torus $T^{2}$, and in terms of circle coordinates $x_{1},x_{2}$ on $T^{2}$ and a suitable transverse coordinate $r$, the contact form $\alpha$ is of Lutz type $\alpha=h_{1}(r)\,\mathrm{d}x_{1}+h_{2}(r)\,\mathrm{d}x_{2}$, and the integral $f$ a function of $r$.
For the case when the integral $f$ is a Morse–Bott function, we also established a neighbourhood theorem for isolated elliptic Reeb orbits in the critical set $\operatorname{Crit}(f)$, that is, an orbit where the Bott integral has a minimum or a maximum; see [6, Theorem 2.4]: In terms of suitable coordinates $(\theta;r,\varphi)$ on a neighbourhood $S^{1}\times D^{2}_{\delta}$ of the critical orbit $S^{1}\times\{0\}$, the Bott integral $f$ equals $f(r)=c\pm r^{2}$, and up to an isotopic deformation of the contact form in a neighbourhood of the critical orbit—changing the Reeb flow but not the Bott integral—the contact form is given by $\mathrm{d}\theta+r^{2}\,\mathrm{d}\varphi$. Also, by Gray stability, the contact structure $\xi=\ker\alpha$ remains unchanged, up to a diffeotopy of $M$, under this deformation.
In the present section we extend the arguments for achieving those results in order to prove neighbourhood theorems for tori and Klein bottles contained in $\operatorname{Crit}(f)$. We then reason as in [6, Section 10.2] to remove those surfaces from the critical set by a $C^{\infty}$-small perturbation of the Bott integral $f$, at the cost of introducing an elliptic and a parallel hyperbolic Reeb orbit into the critical set.
@@ 2.1. Neighbourhood theorems for critical surfaces
Statement (i) follows from the generalised Morse–Bott lemma [3, Lemma 1.7]. In order to prove (ii), we slightly modify the argument for proving [6, Theorem 2.2]. There we constructed covering maps $\mathbb{R}^{2}\rightarrow\{r\}\times T^{2}$ by using the flow of $R$ and $Y$. In the present situation, $Y$ vanishes along the critical set $\{r=0\}$. We circumvent this problem by working with the vector field $Y^{\prime}$ defined by
@@
in other words, we replace $\mathrm{d}f$ in the defining equations (1) for $Y$ by $\mathrm{d}f/r=\pm 2\,\mathrm{d}r$ for $r\neq 0$. This is well defined also in $r=0$, since $r$ is a smooth coordinate on $(-\delta,\delta)$. So $Y^{\prime}$ coincides with $Y/r$ for $r\neq 0$ and extends smoothly as a non-vanishing vector field over $r=0$. The remainder of the argument is now as in the proof of [6, Theorem 2.2]. ∎
Next we describe the neighbourhood of a critical Klein bottle $K$. Recall from [7] or [6, Section 10.1] that a closed tubular neighbourhood $\nu K$ of a Klein bottle $K$ embedded in an orientable $3$-manifold, which is the unique non-trivial interval bundle over $K$, can be described by
@@
The contact form $\alpha$ and the Bott integral $f$ lift to this double cover. In this lift, $\{0\}\times T^{2}=\{r=0\}$ is a component of the critical set. Now, we cannot directly quote the previous result, since we need to have a $\mathbb{Z}_{2}$-equivariant version of that neighbourhood theorem, and we want to achieve a more specific Lutz form. So we need to dig a little deeper into the proof of the neighbourhood theorem for elliptic critical orbits [6, Theorem 2.4], which contains the relevant ideas. At each step in the following argument, we deal with a $\mathbb{Z}_{2}$-equivariant isotopy or a $\mathbb{Z}_{2}$-invariant rescaling of the contact form, all of which are supported in a neighbourhood of $\{r=0\}$. So the actual normal form we finally arrive at is defined on a neighbourhood of $K$ defined by a smaller $\delta$ than the one we started with.
@@
In particular, we have $h_{0}(0)=0$. As in Step 1 of the proof of [6, Theorem 2.4], we first deform $h_{0}$, relative to a neighbourhood of {±$\{\pm\delta\}\times T^{2}$, so that it becomes identically zero near $r=0$; this can be done within the class of odd functions. The convex linear interpolation between the old and the deformed $h_{0}$ gives rise to a strict contact isotopy, defined by a flow along the level surfaces $\{r=\text{const.}\}$. }, and stationary on $\{r=0\}$, just like in the proof of [6, Theorem 2.2]; see equation (3) in that proof and the argument following it.
The $\mathbb{Z}_{2}$-invariance of $\widetilde{R}$, which depends on $r$ only, forces the Reeb vector field to be a multiple of $\partial_{t}$ along $\{r=0\}$. This implies that $h_{1}(0)\neq 0$, and by choosing the signs of the coordinates $r,s$ appropriately, we may assume that $h_{1}(0)>0$ and that the ambient orientation is given by $\mathrm{d}r\wedge\mathrm{d}s\wedge\mathrm{d}\theta$. Then as in Step 2 of the proof of [6, Theorem 2.4], we can make a further deformation such that the contact form looks like $\mathrm{d}s+h_{2}(r)\,\mathrm{d}\theta$ (with a modified, but still odd $h_{2}$) near $\{r=0\}$: for this we need to divide the contact form by an even function $\chi(r)$ equal to $h_{1}(r)$ near $r=0$, and identically equal to $1$ outside a small neighbourhood of $r=0$; so the contact structure remains unchanged. This deformation changes the Reeb vector field, keeping it tangent to the tori $\{r=\text{const.}\}$. }. Thanks to the contact condition, $h_{2}$ satisfies $h_{2}^{\prime}<0$.
@@ 2.2. Removing critical surfaces
For the Klein bottle, we explained in [6, Section 10.2] how to perturb the Bott integral $f$ into one that has two isolated critical Reeb orbits, one elliptic and one hyperbolic, in place of the critical Klein bottle.
@@
In the torus case, we modify the contact form $\alpha=h_{1}(r)\,\mathrm{d}x_{1}+h_{2}(r)\,\mathrm{d}x_{2}$ near $\{r=0\}$ in a way described in the proof of [6, Proposition 2.6]. By a small reparametrisation of the curve $r\mapsto(h_{1}(r),h_{2}(r))$ we may assume that $h_{1}(0)$ and $h_{2}(0)$ are rationally dependent. Then, up to an $\mathrm{SL}(2,\mathbb{Z})$-transformation we may assume that $h_{1}(0)>0$ and $h_{2}(0)=0$, which implies $h_{2}^{\prime}(0)<0$. After a small deformation and a further reparametrisation of the curve $r\mapsto(h_{1}(r),h_{2}(r)$, up to a global constant we may finally assume that $h_{1}$ is identically equal to $1$ near $r=0$, and $h_{2}(r)=-r$ near $r=0$. These modifications change the Reeb vector field, keeping it tangent to the tori $\{r=\text{const.}\}$. }, so $f$ is still a Bott integral. Moreover, near $\{r=0\}$ the new Reeb vector field equals $\partial_{x_{1}}$, so the new Bott integral
@@ 3.1. Euler class and critical orbits
For the case that $L_{f}=\operatorname{Crit}(f)$, that is, in the absence of critical surfaces, the following result was stated implicitly in the proof of [6, Lemma 7.2]. Here and throughout this paper, (co-)homology groups are understood to be with $\mathbb{Z}$-coefficients.
@@ 3.2. The link of critical orbits
In [6, Section 3.2] we gave a summary of the results from [3] concerning the (singular) Liouville foliation defined by the level sets of a Bott integral. The key statement is that up to a diffeomorphism preserving the Liouville foliation, any Bott-integrable Reeb flow can be constructed from some simple building blocks, which are products of a surface (with boundary) with $S^{1}$, or Seifert bundles with singular fibres of order $2$, coming from fixed points of an involution on the surface. The Liouville foliation on these building blocks is defined by the level sets of a Morse–Bott function lifted from a Morse–Bott function on the base, with all fixed points of the involution contained in the critical set of that Morse function.
@@
In [6, Section 3.2.3], we wrote that the Morse–Bott functions on the Seifert fibred building blocks are lifted from Morse functions on the base, when in fact we should have written ”Morse–Bott functions on the base”. The neighbourhood $\nu K$ of a critical Klein bottle is a case in point. As described in detail in [7], $\nu K$ may be thought of as the Seifert bundle over a disc with two singular fibres of order $2$. This corresponds to starting with the annulus $[-\delta,\delta]\times S^{1}$ with Morse–Bott function $f(r,\theta)=r^{2}$ and involution $\tau(r,\theta)=(-r,-\theta)$ having two fixed points. One then obtains $\nu K$ with the described Seifert fibration as the mapping torus of $\tau$.
@@ 4. Seifert manifolds
In [6, Theorem 1.8] we showed that every contact structure on a closed Seifert manifold invariant under the fixed-point free $S^{1}$-action defining the Seifert fibration is Bott integrable. In the present section we deal with overtwisted (but not necessarily $S^{1}$-invariant) contact structures on Seifert manifolds.
Our aim will be to find a Bott-integrable contact structure on any given Seifert manifold that is sufficiently ‘rich’ in the following sense. When we speak of a Bott-integrable contact structure on a manifold with boundary, it is understood that the boundary components are tori, and the contact structure is of Lutz type on collar neighbourhoods of those boundary components. In particular, the boundary is a regular level set of the Bott integral, and the Reeb flow is tangent to the boundary. By the sewing lemma [6, Lemma 3.4], the gluing along such boundaries produces again a Bott-integrable contact structure.
@@
By Eliashberg’s classification [4] of overtwisted contact structures we need only show that every homotopy class of tangent $2$-plane fields on $M$ contains a Bott-integrable overtwisted contact structure. Write $\xi_{0}$ for the given $H_{1}$-complete Bott-integrable contact structure on $M$. We may assume that $\xi_{0}$ is overtwisted; if it is not, we create an elliptic critical orbit (together with a hyperbolic one) as in [6, Section 2.3], and then perform a full Lutz twist, which does not change the homotopy class of $\xi_{0}$ as a tangent $2$-plane field.
Now, given any tangent $2$-plane field $\eta$, there is an obstruction $d^{2}(\xi_{0},\eta)\in H^{2}(M)$ for $\xi_{0}$ to be homotopic to $\eta$ over the $2$-skeleton of $M$. This obstruction is antisymmetric and additive; see [5, Section 4.2]:
@@
see [5, Remark 4.3.4]. Thus, in the presence of $2$-torsion in $H^{2}(M)$ the obstruction class is not detected by the Euler class.
@@
Finally, in order to modify $\xi^{u}$ into a Bott-integrable contact structure homotopic to $\eta$ over all of $M$, it suffices to take a contact connected sum as in [6, Section 6] with a suitable overtwisted contact structure on $S^{3}$, all of which are Bott integrable by [6, Theorem 1.9]. ∎
@@ 4.1.1. Open books
As discussed in [6, Section 7], the manifold $\Sigma\times S^{1}$ may be thought of as an open book with page $\Sigma_{0}$ equal to $\Sigma$ with two open discs removed, and monodromy a couple of Dehn twists (one right-handed, one left-handed) parallel to the two boundary components $\partial\Sigma_{0}\setminus\partial\Sigma$ created by the removal of these discs. For the construction of an $H_{1}$-complete Bott-integrable contact structure on $\Sigma\times S^{1}$, we need the following mild extension of [6, Lemma 3.3]. We formulate it for a single oriented, embedded circle $\gamma\subset\Sigma_{0}$, but it works equally well for a finite collection of pairwise disjoint embedded circles.
@@
Now apply the argument from [6, Lemma 3.3], involving compactly supported cohomology, to the surface obtained by cutting open $\Sigma_{0}$ along $\gamma$ in order to find a primitive $\lambda$ of $\omega$ with the desired properties. ∎
Write $\psi\colon\thinspace\Sigma_{0}\rightarrow\Sigma_{0}$ for the diffeomorphism given by the two boundary-parallel Dehn twists. As shown in [6, Section 7.2], an area form $\omega=\mathrm{d}\lambda$ as just constructed gives rise to a Bott-integrable contact form on $\Sigma\times S^{1}$, where the Bott integral on the mapping torus $\Sigma_{0}(\psi)$ can be taken to be one induced from any Morse function on $\Sigma_{0}$ that only depends on the radial coordinate in the collar neighbourhoods of the boundary (and hence is invariant under $\psi$). On the complement $\Sigma_{0}^{\prime}\subset\Sigma_{0}$ of two collars containing the annuli where the Dehn twists are performed, the mapping torus looks like a product $\Sigma_{0}^{\prime}\times S^{1}$ with contact form $\lambda+C\,\mathrm{d}\theta$, where $C$ is some large positive constant.